# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ Appendix D Gradient flow with reparametrization
Consider a two-layer ReLU network (1) with $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2). Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{M}\subset\operatorname{supp}(\mathcal{B})$. Consider a bijective reparametrization map $G\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}^{3n+1},G(\nu)=\theta$ and the gradient flow in $\nu$-space (119) with $\eta=\Theta(1/n)$. Assume there exists a function $\Phi\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}$ such that $\Phi$ satisfies Assumption 1 and the equality (121) holds for any $\theta=G(\nu)$. Then letting $\nu(\infty)=\lim_{t\rightarrow\infty}\nu(t)$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,G(\nu(\infty)))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to variational problem (7).
@@ E.2 Representation cost for multivariate networks and general initialization
where $\Theta_{\mathrm{ReLU}}^{(d)}=\{\alpha\in C(\operatorname{supp}(\mu))\colon\alpha\text{ is uniformly continuous}\}$. For a given $h\in\mathcal{F}_{\mathrm{ReLU}}^{(d)}$, consider the minimal representation cost problem with the squared $L_{2}$ norm cost:
@@
where $\{\lambda_{i}\}_{i\in[m]}$ are Lagrangian multipliers to be determined. It is clear that either $\bar{\alpha}=0$ or $\bar{\alpha}$ does not have continuous partial derivatives, implying that $\bar{\alpha}p_{\mathcal{W},\mathcal{B}}$ does not lie in $\mathcal{S}(\mathbb{S}^{d-1}\times\mathbb{R})$ even if $p_{\mathcal{W},\mathcal{B}}$ does. A similar observation has been made by Ongie et al. (2020), who note the difficulty in explicitly characterizing the representation cost with the $L_{1}$ norm in the multivariate setting. Some studies have relaxed the Schwartz space condition for the inversion formula of the Radon transform (Helgason, 1999, Theorem 6.2, Ch. I; Jensen, 2004, Theorem 5.1), which recovers a function $f$ on $\mathbb{R}^{d}$ using $\mathcal{R}(f)$. However, to the best of our knowledge, no corresponding relaxation has been established for the inversion formula of the dual Radon transform, which recovers a function $\alpha$ on $\mathbb{S}^{d-1}\times\mathbb{R}$ using $\mathcal{R}^{*}(\alpha)$. The latter is necessary for characterizing the solution to problem (125) with the relation given in (126).
@@ Appendix G Experiment details
Then we optimize $u$ to minimize the above objective function. To solve variational problem (9), we consider two additional sets of constraints (1) $\hat{\mathcal{G}}_{2}(h)=0$; and (2) $\hat{\mathcal{G}}_{3}(h)=0$. This is to ensure $h\in\mathcal{F}_{\mathrm{Abs}}$ according to Proposition 32. Then we optimize $u$ to minimize $\hat{\mathcal{G}}_{1}$. We implement the optimization by CVXPY (Diamond & Boyd, 2016) version 1.4.1. To estimate the $L_{\infty}$-error between the solution to the variational and the trained networks, we evaluate the trained network on $\{t_{i}\}_{i=0}^{500}$, denoted by $\{f_{i}\}_{i=0}^{500}$, and compute the $l_{\infty}$-norm of the difference vector $[f_{0}-h_{0},\ldots,f_{500}-h_{500}]$.
@@ Appendix H Additional numerical experiments
Figure 4: Left: $\ell_{\infty}$-norm of the difference between the final parameter and initial parameter of ReLU networks trained by mirror descent with different unscaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$, plotted against the number of hidden units. Right: spectral norm ($2$-norm) of the difference between the final kernel matrix and the initial kernel matrix from the same training instances as the left panel, plotted against the number of hidden units.
@@
Figure 5: Left panel: $\ell_{\infty}$-norm of the difference between the final parameter and initial parameter of absolute value networks trained by mirror descent with different scaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$, plotted against the number of hidden units. Right panel: spectral norm ($2$-norm) of the difference between the final kernel matrix and initial kernel matrix from the same training instances as the left panel, plotted against the number of hidden units. The inset shows the same data with the y-axis in linear scale instead of log-scale.

### v2

@@ Appendix D Gradient flow with reparametrization
Consider a two-layer ReLU network (1) with one input unit and $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2). Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{M}\subset\operatorname{supp}(p_{\mathcal{B}})$. Consider a bijective reparametrization map $G\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}^{3n+1},G(\nu)=\theta$ and the gradient flow in $\nu$-space (122) with $\eta=\Theta(1/n)$. Assume there exists a function $\Phi\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}$ such that $\Phi$ satisfies Assumption 1 and the equality (124) holds for any $\theta=G(\nu)$. Then letting $\nu(\infty)=\lim_{t\rightarrow\infty}\nu(t)$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,G(\nu(\infty)))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to variational problem (7).
@@ E.2 Representation cost for multivariate networks and general initialization
where $\Theta_{\mathrm{ReLU}}^{(d)}=\{\alpha\in C(\operatorname{supp}(\mu))\colon\alpha\text{ is uniformly continuous}\}$. For a given $h\in\mathcal{F}_{\mathrm{ReLU}}^{(d)}$, consider the minimal representation cost problem with the squared $L^{2}$ norm cost:
@@
where $\{\lambda_{i}\}_{i\in[m]}$ are Lagrangian multipliers to be determined. It is clear that either $\bar{\alpha}=0$ or $\bar{\alpha}$ does not have continuous partial derivatives, implying that $\bar{\alpha}p_{\mathcal{W},\mathcal{B}}$ does not lie in $\mathcal{S}(\mathbb{S}^{d-1}\times\mathbb{R})$ even if $p_{\mathcal{W},\mathcal{B}}$ does. A similar observation has been made by Ongie et al. (2020), who note the difficulty in explicitly characterizing the representation cost with the $L^{1}$ norm in the multivariate setting. Some studies have relaxed the Schwartz space condition for the inversion formula of the Radon transform (Helgason, 1999, Theorem 6.2, Ch. I; Jensen, 2004, Theorem 5.1), which recovers a function $f$ on $\mathbb{R}^{d}$ using $\mathcal{R}(f)$. However, to the best of our knowledge, no corresponding relaxation has been established for the inversion formula of the dual Radon transform, which recovers a function $\alpha$ on $\mathbb{S}^{d-1}\times\mathbb{R}$ using $\mathcal{R}^{*}(\alpha)$. The latter is necessary for characterizing the solution to problem (128) with the relation given in (129).
@@ Appendix G Experiment details
Then we optimize $u$ to minimize the above objective function. To solve variational problem (9), we consider two additional sets of constraints (1) $\hat{\mathcal{G}}_{2}(h)=0$; and (2) $\hat{\mathcal{G}}_{3}(h)=0$. This is to ensure $h\in\mathcal{F}_{\mathrm{Abs}}$ according to Proposition 35. Then we optimize $u$ to minimize $\hat{\mathcal{G}}_{1}$. We implement the optimization by CVXPY (Diamond & Boyd, 2016) version 1.4.1. To estimate the $L^{\infty}$-error between the solution to the variational and the trained networks, we evaluate the trained network on $\{t_{i}\}_{i=0}^{500}$, denoted by $\{f_{i}\}_{i=0}^{500}$, and compute the $\ell_{\infty}$ norm of the difference vector $[f_{0}-h_{0},\ldots,f_{500}-h_{500}]$.
@@ Appendix H Additional numerical experiments
Figure 4: Left: $\ell_{\infty}$ norm of the difference between the final parameter and initial parameter of ReLU networks trained by mirror descent with different unscaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$, plotted against the number of hidden units. Right: spectral norm ($2$-norm) of the difference between the final kernel matrix and the initial kernel matrix from the same training instances as the left panel, plotted against the number of hidden units.
@@
Figure 5: Left panel: $\ell_{\infty}$ norm of the difference between the final parameter and initial parameter of absolute value networks trained by mirror descent with different scaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$, plotted against the number of hidden units. Right panel: spectral norm ($2$-norm) of the difference between the final kernel matrix and initial kernel matrix from the same training instances as the left panel, plotted against the number of hidden units. The inset shows the same data with the y-axis in linear scale instead of log-scale.